Ball 1, with a mass of 100 g and traveling at 14.0 m/s , collides head on with ball 2, which has a mass of 350 g and is initially at rest.

What are the final velocities of each ball if the collision is perfectly elastic?

(vfx)1=  m/s

(vfx)2=  m/s

What are the final velocities of each ball if the collision is perfectly inelastic?

(vfx)1=

m/s

(vfx)2= m/s
Answer #1

here,

mass of 1 , m1 = 0.1 kg

mass of 2 , m2 = 0.35 kg

intial speed of a , u1 = 14 m/s

for perfectly elastic collison

let the final speeds be v1 and v2

using conservation of momentum

m1 * u1 = m1 * v1 + m2 * v2

0.1 * 14 = 0.1 * v1 + 0.35 * v2 ....(1)

and

using conservation of energy

0.5 * m1 * u1^2 = 0.5 * m1 * v1^2 + 0.5 * m2 * v2^2

0.1 * 14^2 = 0.1 * v1^2 + 0.35 * v2^2 ....(2)

from (1) and (2)

v1 = - 7.78 m/s

v2 = 6.22 m/s

for perfectly inelastic collison

let the final speed be v

using conservation of momentum

m1 * u1 = (m1 + m2 )* v

0.1 * 14 = ( 0.1 + 0.35) * v

v = 3.1 m/s

the final speed is 3.1 m/s